OpenStudy (anonymous):

find the derivative of the function y = (1-2sqrt(x))/(1-2xsqrt(x)) find the derivative of the function y = (1-2sqrt(x))/(1-2xsqrt(x)) @Mathematics

OpenStudy (anonymous):

y' = (1-2X^1/2)(1-2XsqrtX) - (1-2x^3/2)(1-2sqrtX) all over (1-2x^3/2)^2

OpenStudy (anonymous):

thats what i got so far

OpenStudy (anonymous):

\[\frac{1-2\sqrt{x}}{1-2x\sqrt{x}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that gives one butt ugly derivative for sure .

OpenStudy (anonymous):

you need all the algebra, or the answer? because even "simplified" it is not that "simple"

OpenStudy (anonymous):

lol i have the answer but i need to know how to get to it

OpenStudy (anonymous):

you need all the algebra, or the answer? because even "simplified" it is not that "simple"

OpenStudy (anonymous):

lorda mercy. looks like you have the first step. \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] now i would like to make life a little simpler by writing \[\frac{2\sqrt{x}-1}{2x\sqrt{x}-1}\] and then put \[f(x)=2\sqrt{x}-1, f'(x)=\frac{1}{\sqrt{x}}\] \[g(x)=2x\sqrt{x}-1,f'(x)=3\sqrt{x}\]

OpenStudy (anonymous):

plugging all this mess in to the "quotient rule" gives \[\frac{(2x\sqrt{x}-1)\times \frac{1}{\sqrt{x}}-(2\sqrt{x}-1)\times 3\sqrt{x}}{(2x\sqrt{x}-1)^2}\]

OpenStudy (anonymous):

now probably best bet is to get rid of the compound fraction by multiplying top and bottom by \[\sqrt{x}\] so give \[\frac{(2x\sqrt{x}-1)-(2\sqrt{x}-1)\times 3x}{\sqrt{x}(2x\sqrt{x}-1)^2}\]

OpenStudy (anonymous):

then some more algebra in the numerator, which i will get you finish because i will probably mess it up

OpenStudy (anonymous):

ok i will try thanks

OpenStudy (anonymous):

you need all the algebra, or the answer? because even "simplified" it is not that "simple"

OpenStudy (anonymous):

yw. just multiply out and it should be good. or course i mean in the numerator. leave the denominator alone. if you have the answer you know what you are working towards, and i think i didn't screw up the first part

OpenStudy (anonymous):

but isnt the derivative of 1-2sqrt(x) = -1/sqrt(x)